OpenStudy (anonymous):

Help Please 3x+6y=42 -7x+8y=-109

OpenStudy (anonymous):

1º devide first euation by 3: x+2y=14 2º multiply it by 7: 7x+14y=98 3º add bouth equations 0x + 22y = - 11 4º solve this for y: y=-1/2 5º substitute this value in x+2y=14 x-1=14 6º solve for x: x=15 so: x=15, y=-1/2